# Simulation-based Approach for Fast Optimal Control of a Stefan Problem with Application to Cell Therapy

### v1

@@ Abstract
This article describes a new, efficient way of finding control and state trajectories in optimal control problems by transformation into a system of differential-algebraic equations (DAEs). The optimal control and state vectors can be obtained via simulation of the resulting DAE system with the selected DAE solver, eliminating the need for an optimization solver. Our simulation-based framework was demonstrated and benchmarked against various optimization-based approaches via four case studies associated with optimization and control of a Stefan problem with application to cell therapy. The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving the same level of accuracy in all cases. The proposed technique offers an efficient and reliable framework for optimal control, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
@@ 1 Introduction
Optimal control of Stefan problems was extensively investigated and proven useful for various industrial applications over the past few decades, with many different objective functions, constraints, and controls (manipulated variables) considered. Some examples are the control of the heating process to satisfy the heating speed and thermoelastic stress constraints by manipulating the furnace temperature (Roubíček, 1986), the stabilization of the moving boundary and temperature/concentration fields by varying the heat flux (Pawlow, 1987), the maximization of the amount of melted solid in the melting process via controlling the heat flux (Silva Neto and White, 1994), the control of the moving boundary to follow the desired path in the solidification process by manipulating the wall temperature (Hinze and Ziegenbalg, 2007), the control of the water level in the drainage basin by varying the discharge velocity (Miyaoka and Kawahara, 2008), the optimization of a thin-film drying process via manipulating the air temperature (Mesbah et al., 2014), and the minimization of the metallurgical length (ML) deviation in steel casting via controlling the boundary heat flux (Chen et al., 2019).
@@
All the aforementioned studies rely on optimization algorithms/solvers to obtain the optimal control trajectories. The widely used approach is to replace the time-varying control vector by a finite-dimensional parameterization (e.g., a spline) and carry out numerical discretization to transform the dynamic optimization into a nonlinear algebraic program (Kishida et al., 2013; Scott et al., 2018; Nolasco et al., 2021, for example). This overall approach has many variations, including methods that sequentially switch between solving the numerical discretization of the underlying partial differential-algebraic equations (PDAEs) and running an algebraic optimizer, or simultaneously solving a single large sparse optimization with the numerical discretization equations as explicit constraints (Neuman and Sen, 1974; Tsang et al., 1975; Mellefont and Sargent, 1978; Sargent and Sullivan, 1978; Biegler, 2007; Nolasco et al., 2021). Numerous optimal control algorithms have been developed for efficient solutions to facilitate real-time applications such as model predictive control (e.g., see detailed discussions by Rodrigues et al. (2014) and Nolasco et al. (2021)). Alternatively, Berliner et al. (2022) showed that it is possible to reformulate some optimal control problems as a mixed continuous-discrete system of index-1 PDAEs. The PDAE system is then solved numerically by feeding the differential-algebraic equations (DAEs) obtained by spatial discretization into a DAE solver. In this approach, the optimal control vector is obtained directly by a DAE solver, without using any optimization solver, resulting in a highly computationally efficient solution to the optimal control problem. Recently, this approach was used to solve optimal control problems associated with a Stefan problem (Srisuma et al., 2024).
@@
Although optimal control with Stefan problems has been explored for a wide range of processes, applications to cell therapy have not been available. Previous studies showed that accurate prediction, control, and optimization of cryopreservation and cell thawing can improve the viability and quality of the resulting cells, which directly benefits cell therapy (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Hunt, 2019; Bojic et al., 2021; Cottle et al., 2022; Uhrig et al., 2022). These benefits and case studies therefore motivate the development of an efficient Stefan problem-based optimal control algorithm for cell therapy.
@@
    derive the simulation-based approach for optimal control of a Stefan problem for the cell thawing process in cell therapy;
@@
    generalize the simulation-based approach for both index-1 and high-index DAE systems;
@@
    demonstrate and benchmark the simulation-based algorithm against various optimization-based optimal control algorithms; and
@@
    apply the proposed approach to efficiently and accurately solve several practical problems associated with optimization and control of cell thawing.
@@ 2 Stefan Problem and Cell Thawing
The system used for the demonstration of our optimal control technique in this article concerns cell thawing, which is a process used in cell therapy before cells are introduced to the patients (Fig.  1). During thawing, energy is continuously supplied by a heater to thaw the material in a vial. A mechanistic model of cell thawing can be formulated as a Stefan problem (Srisuma et al., 2023b). Heat transfer in the solid and liquid domains can be described by the energy balance equation,
@@
In dimensionless form, the model consists of three main parts which describe the evolution of the solid temperature (4), the liquid temperature (5), and the interface position (6), respectively. The energy is continuously supplied to the system by the heater with the temperature $\Theta_{b}$. The system initially consists of pure solid, which corresponds to $S=1$; i.e., the interface position is at the outer boundary. As the thawing progresses, the interface position decreases and eventually becomes 0, indicating complete melting or pure liquid. The maximum temperature (dimensionless) is equal to 1, which is equivalent to $37\,^{\circ}$, the maximum temperature recommended for cell thawing (Baboo et al., 2019; Uhrig et al., 2022). The minimum temperature (dimensionless) is equal to 0, corresponding to the melting point of $-2\,^{\circ}$.
@@
In this case, only the liquid temperature $\Theta_{2}$ is considered because the solid temperature $\Theta_{1}$ is equal to the melting point, which does not change with time. Note that, as the problem is defined in the radial direction, the average temperature over the cross section could also be used instead of (10) for more accurate results. In any case, the implementation and algorithms presented later in this article are still valid for both definitions.
@@ 3.1 Optimal control formulation
The control (aka decision variable, manipulated variable, input) of this optimal control problem is the heater temperature $\Theta_{b}(\tau)$. The mechanistic model (4)–(9) describes the physics of cell thawing, and thus serves as constraints. The lower and upper bounds of the heater temperature are represented by (12).
@@ 3.2 Optimization-based approach
A detailed discussion on numerical algorithms to solve optimal control problems can be found in Nolasco et al. (2021). The common technique for solving optimal control problems is to discretize the partial differential equation (PDE) and ordinary differential equation (ODE) constraints, parameterize the time-varying control vector, and solve the resulting optimization problem numerically with a proper optimizer. This approach relies on the convergence of the optimization algorithm/solver to obtain the optimal control trajectory, and hence we denote this approach as the optimization-based approach. In this article, six different optimization-based implementations that have been used widely in optimization and optimal control applications, including on a variety of software tools and programming languages, are considered:
  • opt_Ipopt: The control vector is parameterized, and the resulting optimization is numerically solved by IPOPT (Wächter and Biegler, 2006), an open-source optimization solver for large-scale nonlinear optimization that has been implemented in various nonlinear and optimal control software packages. The IPOPT solver is employed in MATLAB via the OPTI Toolbox (Currie and Wilson, 2012). The system of ODEs is integrated by ode15s.
  • opt_fmincon: The control vector is parameterized, and the resulting optimization is numerically solved by fmincon, a built-in optimization solver for constrained nonlinear optimization in MATLAB. The system of ODEs is integrated by ode15s.
@@ 3.3 Simulation-based approach
The approach in this article is inspired by a recent article on the optimal control of lithium-ion batteries (Berliner et al., 2022). The approach was motivated by the observation that the optimal control trajectories for that application moved from one active inequality constraint to another over time. That information was used to transform the optimal control problem into a mixed continuous-discrete system of index-1 DAEs in which existing software could be used to transition between active inequality constraints. This simulation-based technique eliminated the need for an optimization solver, resulting in a much more computationally efficient solution. No control vector parameterization is needed.
@@ 3.3.1 Solution of high-index DAEs
In this work, we generalize the simulation-based technique for high-index DAE systems and apply the approach to solve multiple cell thawing problems. Various numerical algorithms for solving index-1 DAEs, including the implementation used in Berliner et al. (2022), fail when the differential index is higher than 1 (Petzold, 1982), and an increase in the differential index of DAEs leads to more difficulties in obtaining the numerical solution to DAEs (Campbell et al., 2008). Many current DAE solvers are only capable of solving index-1 DAEs, so it is usually recommended to perform index reduction (e.g., reduceDAEIndex in MATLAB and dae_index_lowering in Julia) to transform a high-index system into an equivalent index-1 system, and then solve the resulting index-1 DAEs (Shampine et al., 1999; Campbell et al., 2008). This technique is usually suggested because index-1 DAEs can be solved easily by most DAE solvers, and so does not require specialized high-index DAE solvers. Nevertheless, there are two major drawbacks associated with index reduction. First, the index reduction process is costly when the number of states and differential index are huge. Second, index reduction could introduce a large number of new variables (sometimes called dummy derivatives (Mattsson and Söderlind, 1993)) to replace high-order derivatives, which unnecessarily increases the complexity of the problem. These issues are not commonly discussed in the literature as they become significant when the differential index is sufficiently large, e.g., Problem 3 in our case studies. Thus, this approach does not match our objective of accelerating the optimal control algorithm.
An alternative to the index reduction approach is to fully discretize a system of high-index DAEs (e.g., using a collocation method) and solve the discretized equations (Campbell et al., 2008). The specialized DAE solver in GEKKO employs the method of orthogonal collocation on finite elements to solve high-index DAEs (Beal et al., 2018). The main advantage of this implementation is that it does not have any constraint or limit on the differential index because index reduction and differentiation are not required (e.g., see detailed discussions by Hedengren et al. (2014) and Beal et al. (2018)). Hence, our DAE solutions for high-index systems rely on the DAE solver implemented in GEKKO.
@@ 3.3.2 Proposed algorithm
Second, treat the control variable ($\Theta_{b}$ in this case) as a state instead of a decision variable. A consistent initial condition is required for this new state, and this initialization can be done in several ways, depending on the solver. For example, in GEKKO, the DAE solver does not strictly require a consistent initial condition, and so we can first guess some random initial condition and then run the DAE solver once to find the correct initial condition. Another possibility is to solve the optimization locally at the beginning of the process. Either approach requires minimal effort and computation, and so does not impact the overall computational performance. The initial condition for this new state is denoted by
@@
After obtaining the DAEs and initial conditions, the final step is to solve the resulting DAEs (13) and (14). As the control is now treated as a state, the optimal control vector can be obtained directly from the DAE solver.
This DAE reformulation technique inherently assumes that at least one of the constraints, bounds, or algebraic equations resulting from reformulating the objective function is active. For many practical problems in which this strong condition does not hold, the simulation-based approach is not guaranteed to produce an optimal solution, which is discussed in more detail in Section 4.3. Changes in the active constraints or bounds can be incorporated in the simulation using a simulation-switching technique, with Section 4.4 providing a discussion that includes references to several examples.
@@ 3.4 Implementation
Both optimization- and simulation-based techniques were implemented in MATLAB R2023a following the procedures introduced in Sections 3.2 and 3.3, with GEKKO called and executed in Python 3.10. The simulations were performed on a computer equipped with an Intel® Core™ i9-13950HX Processor CPU (24 cores) and 128 GB RAM on 64-bit Windows 11. The choice and justification of important solver options are given in Appendix B.
@@ 4 Case Studies
Figure 2: Trajectories for the optimal (A) heater temperature $\Theta_{b}$, (B) interface position $S$, (C) interface velocity $dS/d\tau$, (D) average temperature $\Theta_{\textrm{avg}}$, and (E) rate of temperature change $d\Theta_{\textrm{avg}}/d\tau$ in Problem 1.
@@ 4.1.1 Problem description and formulation
This first problem is simple and intuitive, with the solution known a priori, to demonstrate the simulation-based technique and validate all of our algorithms.
In general, it has been suggested that the heating process should be done rapidly to avoid potential damages to the cells and maintain high viability (Terry et al., 2010; Baboo et al., 2019; Hunt, 2019; Uhrig et al., 2022), which is equivalent to the optimal control problem
@@
The solution to this optimal control problem is intuitive. To minimize the thawing time, the heater temperature should be fixed at its upper bound to provide the maximum heating power throughout the process.
@@ 4.1.3 Solution comparison
Every solution method gives the same optimal solution (Fig.  2); the heater temperature $\Theta_{b}$ is at its upper bound of 1 all the time, ensuring that the algorithms and implementation are correct. The wall-clock time required for the simulation-based technique (sim_DAE) is more than an order of magnitude lower than every optimization-based approach (Table 1). Optimization with GEKKO is the slowest method, followed by IPOPT, CasADi and fmincon. GEKKO’s ODE solvers cannot perform adaptive time-stepping, which could lead to significantly slower computation. Parallel computing fails to accelerate fmincon. The simulation-based approach is the most computationally efficient algorithm in this case study.

### v2

@@ Abstract
This article describes a new, efficient way of finding control and state trajectories in optimal control problems by reformulation as a system of differential-algebraic equations (DAEs). The optimal control and state vectors can be obtained via simulation of the resulting DAE system with the selected DAE solver, eliminating the need for an optimization solver. Our simulation-based approach is demonstrated and benchmarked against various optimization-based algorithms via four case studies associated with the optimization and control of a Stefan problem for cell therapy. The simulation-based approach is faster than every optimization-based method by more than an order of magnitude while giving similar/better accuracy in all cases. The solution obtained from the simulation-based approach is guaranteed to be optimal provided that at least one constraint or algebraic equation resulting from the reformulation remains active at all times. The proposed technique offers an efficient and reliable framework for optimal control, serving as a promising alternative to the traditional techniques in applications where speed is crucial, e.g., real-time online model predictive control.
@@ 1 Introduction
Optimal control of Stefan problems has been extensively investigated and proven useful for various industrial applications over the past few decades, with many different objective functions, constraints, and controls (manipulated variables) considered. Some examples are the control of a heating process to satisfy constraints on the heating speed and thermoelastic stress by manipulating the furnace temperature (Roubíček, 1986), the stabilization of the moving boundary and temperature/concentration fields by varying the heat flux (Pawlow, 1987), the maximization of the amount of melted solid in a melting process via controlling the heat flux (Silva Neto and White, 1994), the control of the moving boundary to follow the desired path in a solidification process by manipulating the wall temperature (Hinze and Ziegenbalg, 2007), the control of the water level in a drainage basin by varying the discharge velocity (Miyaoka and Kawahara, 2008), the optimization of a thin-film drying process via manipulating the air temperature (Mesbah et al., 2014), and the minimization of the metallurgical length deviation in steel casting by controlling the boundary heat flux (Chen et al., 2019).
@@
All the aforementioned studies rely on optimization algorithms/solvers to obtain the optimal control trajectories. The widely used approach is to replace the time-varying control vector by a finite-dimensional parameterization (e.g., a spline) and carry out numerical discretization to transform the dynamic optimization into a nonlinear algebraic program (Kishida et al., 2013; Scott et al., 2018; Nolasco et al., 2021, for example). This overall approach has many variations, including methods that sequentially switch between solving the numerical discretization of the underlying partial differential-algebraic equations (PDAEs) and running an algebraic optimizer, or simultaneously solving a single large sparse optimization with the numerical discretization equations as explicit constraints (Neuman and Sen, 1974; Tsang et al., 1975; Mellefont and Sargent, 1978; Sargent and Sullivan, 1978; Biegler, 2007; Nolasco et al., 2021). Numerous optimal control algorithms have been developed for efficient solutions to facilitate real-time applications such as model predictive control (e.g., see detailed discussions by Rodrigues et al. (2014) and Nolasco et al. (2021)). Alternatively, Berliner et al. (2022) showed that it is possible to reformulate some optimal control problems as a mixed continuous-discrete system of index-1 PDAEs. The PDAE system is then solved numerically by feeding the differential-algebraic equations (DAEs) obtained by spatial discretization into a DAE solver. In this approach, the optimal control vector is obtained directly from the DAE solver, without using any optimization solver, resulting in a highly computationally efficient solution to the optimal control problem. Recently, this approach was used in the optimal control of a Stefan problem (Srisuma et al., 2024).
@@
Although optimal control of Stefan problems has been explored for a wide range of processes, applications to cell therapy are lacking. Previous studies have shown that accurate prediction, control, and optimization of cryopreservation and cell thawing can improve the viability and quality of the resulting cells, which directly benefits cell therapy (Seki and Mazur, 2008; Jang et al., 2017; Baboo et al., 2019; Hunt, 2019; Bojic et al., 2021; Cottle et al., 2022; Uhrig et al., 2022). These benefits and case studies therefore motivate the development of an efficient Stefan problem-based optimal control algorithm for cell therapy.
@@
    derive the simulation-based approach for optimal control of a Stefan problem applied to the cell thawing process in cell therapy;
@@
    generalize the simulation-based approach to both index-1 and high-index DAE systems;
@@
    demonstrate and benchmark the simulation-based algorithm against various optimization-based optimal control algorithms;
@@
    apply the proposed approach to efficiently and accurately solve several practical problems associated with the optimization and control of cell thawing; and
  5. 5.
    discuss its limitation and condition required for the approach to provide an optimal solution.
@@ 2 Stefan Problem and Cell Thawing
The system used for the demonstration of our optimal control technique in this article concerns cell thawing, which is a process used in cell therapy before cells are introduced into the patients (Fig. 1). During thawing, energy is continuously supplied by a heater to thaw the material in a vial. A mechanistic model of cell thawing can be formulated as a Stefan problem (Srisuma et al., 2023b). Heat transfer in the solid and liquid domains can be described by the energy balance equation,
@@
In dimensionless form, the model consists of three main parts that describe the evolution of the solid temperature (4), the liquid temperature (5), and the interface position (6), respectively. The system is initially in the solid state, which corresponds to $S=1$; i.e., the interface position is at the outer boundary. As the thawing progresses, the interface position decreases and eventually becomes 0, indicating complete melting. The heater temperature is $\Theta_{b}$, which can be fixed or varied with time. The maximum dimensionless temperature is 1, which is equivalent to $37\,^{\circ}$, the highest temperature recommended for cell thawing (Baboo et al., 2019; Uhrig et al., 2022). The minimum dimensionless temperature is 0, corresponding to the melting point of $-2\,^{\circ}$.
@@
In this case, only the liquid temperature $\Theta_{2}$ is considered because the solid temperature $\Theta_{1}$ is equal to the melting point, which does not change with time. Note that, as our problem is defined in the radial direction, the average temperature over the cross section could also be used instead of (10) for more accurate results. In any case, the implementation and algorithms presented later in this article are still valid for both definitions.
@@ 3.1 Optimal control formulation
The control (aka decision variable, manipulated variable, input) of this optimal control problem is the heater temperature $\Theta_{b}(\tau)$. The mechanistic model (4)–(9) describes the physics of cell thawing and thus serves as the constraints of the optimization. The lower and upper bounds on the heater temperature are represented by (12).
@@ 3.2 Optimization-based approach
A detailed discussion on numerical algorithms for solving optimal control problems can be found in Nolasco et al. (2021). The common technique for solving optimal control problems is to discretize the partial differential equation (PDE) and ordinary differential equation (ODE) constraints, parameterize the time-varying control vector, and solve the resulting optimization numerically with a proper optimizer. This approach relies on the convergence of the optimization algorithm/solver to obtain the optimal control trajectory, and hence we denote this approach as the optimization-based approach. In this article, six different optimization-based implementations that have been used widely in optimization and optimal control applications, including a variety of software tools and programming languages, are considered:
  • opt_Ipopt: The control vector is parameterized, and the resulting optimization is numerically solved by IPOPT (Wächter and Biegler, 2006), an open-source optimization solver for large-scale nonlinear optimization that has been implemented in various nonlinear and optimal control software packages. The IPOPT solver is employed in MATLAB via the OPTI Toolbox (Currie and Wilson, 2012). The system of ODEs is integrated using ode15s.
  • opt_fmincon: The control vector is parameterized, and the resulting optimization is numerically solved by fmincon, a built-in optimization solver for constrained nonlinear optimization in MATLAB. The system of ODEs is integrated using ode15s.
@@ 3.3 Simulation-based approach
The approach in this article is inspired by a recent article on the optimal control of lithium-ion batteries (Berliner et al., 2022). The approach was motivated by the observation that the optimal control trajectories for that application moved from one active inequality constraint to another over time. That information was used to transform the optimal control problem into a mixed continuous-discrete system of index-1 DAEs, in which existing software can be used to transition between active inequality constraints. This simulation-based technique eliminates the need for an optimization solver, resulting in a much more computationally efficient solution. No control vector parameterization is needed.
@@ 3.3.1 Solution of high-index DAEs
In this work, we extend the simulation-based technique to handle high-index DAE systems and apply the approach to solve multiple cell thawing problems. Various numerical algorithms for solving index-1 DAEs, including the implementation used in Berliner et al. (2022), fail when the differential index is higher than 1 because an increase in the differential index of DAEs introduces more difficulties in obtaining numerical solutions (Petzold, 1982; Campbell et al., 2008). Many current DAE solvers are only capable of solving index-1 DAEs, so it is usually recommended to perform index reduction (e.g., reduceDAEIndex in MATLAB and dae_index_lowering in Julia) to transform a high-index system into an equivalent index-1 system, and then solve the resulting index-1 DAEs (Shampine et al., 1999; Campbell et al., 2008). This technique is generally preferable because index-1 DAEs can be solved easily by most DAE solvers, eliminating the need for specialized high-index DAE solvers. Nevertheless, there are two major drawbacks associated with index reduction. First, the index reduction process is costly when the number of states and differential index are huge. Second, index reduction could introduce a large number of new variables, sometimes called dummy derivatives (Mattsson and Söderlind, 1993), to replace high-order derivatives, which unnecessarily increases the complexity of the problem. These issues are not commonly discussed in the literature as they become significant when the differential index is sufficiently large, e.g., Problem 3 in our case studies. Thus, this approach does not match our objective of accelerating the optimal control algorithm.
An alternative to the index reduction approach is to fully discretize a system of high-index DAEs (e.g., using a collocation method) and solve the discretized equations (Campbell et al., 2008). The specialized DAE solver in GEKKO employs the method of orthogonal collocation on finite elements to solve high-index DAEs (Beal et al., 2018). The main advantage of this implementation is that it does not entail any index reduction or dummy derivatives, making it more efficient when dealing with some high-index problems (e.g., see detailed discussions by Hedengren et al. (2014) and Beal et al. (2018)). Hence, our DAE solutions for high-index systems rely on the DAE solver implemented in GEKKO.
@@ 3.3.2 Proposed algorithm
Second, treat the control variable ($\Theta_{b}$ in this case) as a state instead of a decision variable. A consistent initial condition is required for this new state, and the initialization can be done in several ways, depending on the solver. For example, in GEKKO, the DAE solver does not strictly require a consistent initial condition, so it is possible to start with some reasonable guess and then run the DAE solver once to find a more accurate initial condition. Another possibility is to solve the optimization locally at the beginning of the process. Either approach requires minimal effort and computation, and therefore does not impact the overall computational performance. The initial condition for this new state is denoted by
@@
After obtaining the DAEs and initial conditions, the final step is to solve the resulting DAEs (13) and (14). As the control is now treated as a state, the optimal control vector is obtained automatically from the DAE solver.
This DAE reformulation technique inherently assumes that at least one of the inequality constraints, bounds, or algebraic equations resulting from reformulating the objective function is active. For problems in which this condition does not hold, the simulation-based approach is not guaranteed to produce an optimal solution, which is discussed in more detail in Section 4.3. Changes in the active constraints or bounds can be incorporated in the simulation using a switching technique, with Section 4.4 providing a discussion that includes references to several examples.
@@ 3.4 Implementation
In this work, both optimization- and simulation-based techniques were implemented in MATLAB R2023a following the procedures introduced in Sections 3.2 and 3.3, with GEKKO called and executed in Python 3.10. All simulations were performed on a computer equipped with an Intel® Core™ i9-13950HX processor (24 cores) and 128 GB RAM running 64-bit Windows 11. The choice and justification of important solver options are given in Appendix A.
@@ 4 Case Studies
Figure 2: Trajectories of the optimal (A) heater temperature $\Theta_{b}$, (B) interface position $S$, (C) interface velocity $dS/d\tau$, (D) average temperature $\Theta_{\textrm{avg}}$, and (E) rate of temperature change $d\Theta_{\textrm{avg}}/d\tau$ for Problem 1.
@@ 4.1.1 Problem description and formulation
This first problem is simple and intuitive, with the solution known a priori, to demonstrate the simulation-based technique and validate all algorithms.
In general, it has been suggested that the heating process should be done rapidly to avoid potential damage to the cells and maintain high viability (Terry et al., 2010; Baboo et al., 2019; Hunt, 2019; Uhrig et al., 2022), which is equivalent to the optimal control problem
@@
The solution to this optimal control problem is straightforward. To minimize the thawing time, the heater temperature should be fixed at its upper bound to deliver the maximum heating power throughout the process.
@@ 4.1.3 Solution comparison
Every solution method gives the same optimal solution (Fig. 2); the heater temperature $\Theta_{b}$ is at its upper bound throughout the process, indicating that our algorithms and implementations are correct. The wall-clock time required for the simulation-based technique (sim_DAE) is several orders of magnitude lower than those of the optimization-based approaches (Table 1). Optimization with GEKKO is the slowest method, followed by IPOPT, CasADi and fmincon. GEKKO’s ODE solvers cannot perform adaptive time-stepping, which could lead to significantly slower computation. Parallel computing fails to accelerate fmincon for this simple problem. Overall, the simulation-based approach is the most computationally efficient algorithm.